OpenStudy (anonymous):

Find the value of 21 + 4(32 - 5). 25 37 100

OpenStudy (muscrat123):

hi again

OpenStudy (anonymous):

haha hey

OpenStudy (muscrat123):

no other options?

OpenStudy (calculusxy):

PEMDAS

OpenStudy (anonymous):

no but its not 32 its \[3^{2}\]

OpenStudy (muscrat123):

oh ok

OpenStudy (calculusxy):

\[21 + 4(3^2 - 5)\]

OpenStudy (muscrat123):

it is c

OpenStudy (calculusxy):

Is that the expression?

OpenStudy (anonymous):

yea @calculusxy

OpenStudy (muscrat123):

may i explain? @audieeeuniverse

OpenStudy (anonymous):

yes @muscrat123

OpenStudy (muscrat123):

first u do the ( ). do u know how to do exponents?

OpenStudy (anonymous):

yea

OpenStudy (muscrat123):

what is 3 x 3?

OpenStudy (anonymous):

9

OpenStudy (muscrat123):

9 - 5 =

OpenStudy (anonymous):

3

OpenStudy (muscrat123):

nope

OpenStudy (anonymous):

oh wait 4 duh

OpenStudy (muscrat123):

lol. so the ( ) have a total of 4

OpenStudy (anonymous):

okay

OpenStudy (muscrat123):

and 21+ 4 = 25

OpenStudy (anonymous):

okay

OpenStudy (calculusxy):

@muscrat123 That's not correct

OpenStudy (calculusxy):

4 is with the parenthesis so you need to multiply: \[4(4)\]

OpenStudy (muscrat123):

oh yeah1! srry it looked like it was 21 + 4. ill fix this

OpenStudy (calculusxy):

So now what you have is: 21 + 4(4) 21 + 16

OpenStudy (calculusxy):

Add those up to get your answer

OpenStudy (muscrat123):

@audieeeuniverse solve 21 + 16 = ur answer

OpenStudy (muscrat123):

which is B

OpenStudy (anonymous):

37? yea thats b

OpenStudy (muscrat123):

YAY!!!!!

OpenStudy (anonymous):

thank you all

OpenStudy (muscrat123):

the 4 is with the ( ), right?

OpenStudy (muscrat123):

and np. anytime

OpenStudy (anonymous):

you the 4 it right in from of them

OpenStudy (anonymous):

*yea

OpenStudy (muscrat123):

yep! ok. any other ?s

OpenStudy (anonymous):

yea ive got tons of questions

OpenStudy (muscrat123):

i can help!

OpenStudy (muscrat123):

just message me

OpenStudy (anonymous):

okay thanks, do you want me to make a new thing and tag you so you can get medals?

OpenStudy (muscrat123):

actually, yes that would b better. THX!!!!!!

OpenStudy (anonymous):

haha okay thank you